OpenStudy (agray99):

please help me solve the equation 3/2 (x-2)-5=19

OpenStudy (agray99):

and 6=1/5w+7/5w-4

OpenStudy (igreen):

\(\dfrac{3}{2}(x - 2) - 5 = 19\) Is this it?

OpenStudy (igreen):

@agray99

OpenStudy (agray99):

i can't veiw your answer

OpenStudy (igreen):

I didn't give an answer..lol.

OpenStudy (igreen):

Just asking if that's the equation..

OpenStudy (agray99):

the answer to number 1 is x=18 and number 2 is w=25/4 but i need to find the processs

OpenStudy (igreen):

Okay, well first we distribute 3/2 into the parenthesis: \(\dfrac{3}{2}(x - 2)\) To do this we multiply it to both 'x' and '-2'. \(\dfrac{3}{2} \times x = \dfrac{3}{2}x\) \(\dfrac{3}{2} \times -2 = -3\) So we have: \(x - 3 - 5 = 19\) Can you finish that? @agray99

OpenStudy (igreen):

Sorry we have: \(\dfrac{3}{2}x - 3 - 5 = 19\)

OpenStudy (igreen):

Can you finish that? @agray99 Just subtract -3 - 5, add that to both sides, then multiply 2/3 to both sides.

OpenStudy (agray99):

i'm confused as to how we got all that

OpenStudy (igreen):

We distribute 3/2 into the parenthesis..the parenthesis has 'x' and -2' inside of it..distributing 3/2 will be multiplying 3/2 to 'x' and '-2'..and when I did that I came up with 3/2x - 3